OpenStudy (anonymous):

What is the value of h? 4ʰ=1/64

Nnesha (nnesha):

can you write 64 in terms of base 4 ?

OpenStudy (anonymous):

Yes

Nnesha (nnesha):

okay so 4 to what power = 64 ?

OpenStudy (anonymous):

16

Nnesha (nnesha):

4^{16} = 64 ?

OpenStudy (anonymous):

XD im stupid sorry its 3

Nnesha (nnesha):

yes right so \[\huge\rm 4^h =\frac{ 1 }{ 4^3 }\] so we can write 64 as 4^3 now move the 4^3 to the top(numerator) flip the fraction when you flip the fraction sign of the `exponent` would change

OpenStudy (anonymous):

So it would be 4^-3?

Nnesha (nnesha):

yes right \[\huge\rm 4^h=4^{-2}\] bases are the same so you can cancel thm out

Nnesha (nnesha):

here is an exxample \[\huge\rm \cancel{2}^x=\cancel{2}^y~~~~~x=y\]

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

So H = -3

Nnesha (nnesha):

good